# Think Together and Work Better: Combining Humans’ and LLMs’ Think-Aloud Outcomes for Effective Text Evaluation

### v1

@@ 4.4. Qualitative analysis
We conducted a focused group interview (FGI) (Morgan, 1996; Dilshad and Latif, 2013) with four human experts (E1, E2, E3, and E4) to qualitatively analyze the attributes produced by humans (A) and LLMs (B). The experts compared both sets of attributes by responding to open-ended questions for each dimension (see Appendix G for the details of questions). These questions allowed us to gain deeper insights into the attributes and encouraged detailed expert feedback. Three of the experts had participated in the TA, while one had not. Initially, we planned for all four TA participants to conduct the qualitative analysis, assuming their familiarity would lead to more refined insights. However, one expert with linguistics expertise withdrew for a personal reason, so we invited another linguist, who had been also researching large language models, to participate. After collecting the experts’ responses through email, we identified keywords mentioned by the majority (at least three experts). We then filtered out answers containing the keywords. The results will be further discussed in section 5.3.2 and 5.3.3.
@@ 5. Results
5.1. Performance comparison with baselines (RQ1)
@@
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
@@
ROUGE-L 0.172 0.122 0.240 0.167 0.317 0.235 0.420 0.301 0.260 0.187
BLEU 0.028 0.019 -0.079 -0.057 0.036 0.030 0.423 0.308 0.023 0.017
METEOR 0.009 0.008 0.011 0.008 0.101 0.074 0.275 0.182 0.055 0.040
BERTScore -0.219 -0.151 -0.156 -0.113 -0.109 -0.076 0.462 0.476 -0.188 -0.134
MOVERScore 0.340 0.252 0.345 0.247 0.384 0.277 -0.434 -0.315 0.349 0.268
BARTScore 0.475 0.343 0.395 0.285 0.457 0.341 0.602 0.455 0.451 0.330
UniEval 0.567 0.411 0.605 0.447 0.613 0.476 0.640 0.476 0.555 0.412
@@
G-Eval 0.412 0.393 0.354 0.333 0.338 0.322 0.361 0.321 0.366 0.342
CheckEval 0.440 0.347 0.393 0.327 0.600 0.515 0.305 0.252 0.436 0.360
InteractEval 0.583 0.536 0.630 0.604 0.734 0.702 0.614 0.576 0.640 0.605
@@
G-Eval 0.619 0.470 0.629 0.482 0.664 0.517 0.617 0.464 0.632 0.483
CheckEval 0.573 0.428 0.632 0.493 0.706 0.611 0.570 0.438 0.620 0.493
InteractEval 0.650 0.603 0.798 0.768 0.783 0.760 0.626 0.587 0.668 0.640
Table 2. Sample-level correlations of different dimensions on SummEval benchmark dataset measured by Spearman ($\rho$) and Kendall tau ($\tau$). The best results are highlighted in bold, and the second-best results are underlined.
To evaluate the performance of InteractEval, we conduct a comparison against baseline evaluation methods. Building on previous studies (Fu et al., 2023; Fabbri et al., 2021; Lee et al., 2024b; Liu et al., 2023), we analyze the correlation between InteractEval and human annotation scores using Spearman and Kendall-tau coefficients to validate the reliability of our evaluations. For a fair comparison, we use the performance data of the non-LLM-based traditional method, G-Eval (GPT-4) and CheckEval as reported in (Lee et al., 2024b), since their experiments were conducted using the same experimental settings (sample-level correlation). We also reproduced the correlation results of G-Eval (GPT-3.5-Turbo) at the sample level using the source code provided by the original authors  [^8] https://github.com/nlpyang/geval.git, as none of previous work has included the results.
@@
Table 2 presents the main results, classifying the four dimensions into two key features by which generated texts are typically assessed: internal quality and external alignment (Zhu and Bhat, 2020; Wang et al., 2024c). InteractEval with GPT-4 outperforms the traditional non-LLM-based metrics, such as ROUGE-L, BLEU, METEOR, BERTScore, MOVERScore, and BARTScore, in terms of correlation coefficients. Except for the Spearman correlation in the ”Relevance” dimension, InteractEval with GPT-4 generally outperforms UniEval. As noted in (Lee et al., 2024b), this suggests that using a checklist enables an LLM to evaluate generated text in a manner more aligned with human judgment than traditional automatic metrics.
@@
When compared to G-Eval and CheckEval with the identical version of GPT, our model demonstrates superior performance. Moreover, even with GPT-3.5-Turbo, our framework outperforms other LLM-based baselines with GPT-4 or shows comparable performance. This indicates the cost-saving potential while ensuring reasonable performance of our proposed framework (Arefeen et al., 2024; Renze and Guven, 2024; Shekhar et al., 2024). These results highlight the broad capability of our proposed framework, showing that using TA and the checklist constructed through the combinatory TA of humans and LLMs lead to accurate and cost-effective evaluations across various text quality dimensions. Since InteractEval performs best when using GPT-4 as the evaluator, the remaining experiments were conducted with GPT-4.
5.2. Analysis of TA effectiveness (RQ2)
@@
Table 3. Sample-level correlations of different dimensions on the SummEval benchmark dataset were measured using Spearman ($\rho$) and Kendall-tau ($\tau$). Entire models utilized GPT-4 as the evaluator. The model names in parentheses in InteractEval refer to the models used for the TA process. The best results are highlighted in bold, and the second-best results are underlined, presented evaluator-wise.
@@
The proposed framework initially collects various attributes from humans and LLMs that should be considered while designing checklists per dimension. Unlike previous methods, we apply TA to encourage LLMs to generate a wider range of text attributes, which form the basis for checklists across different evaluation dimensions. To examine how the TA process affects the evaluation performance of our framework, we compare its performance when using checklists constructed from a single LLM’s attributes under the TA process against other LLM-based baselines. The performance is shown in Table 3. As a result, the performance of InteractEval across four dimensions varies depending on the LLM used for the TA process. Still, our method with the single-LLM-TA outperforms other LLM-based baselines including the average, except in Spearman correlation for Relevance. Nevertheless, InteractEval using the outcomes of single-LLM TA with GPT-4 surpasses the baselines on average.
These results, shown in Table 3, suggest that even with just one LLM, the TA process leads to more refined checklists, improving evaluation outcomes. By encouraging divergent thinking (Ericsson and Simon, 1980; Ericsson, 6; Lewis, 1982), the TA process enables LLMs to capture more detailed aspects of the dimension rubrics based on sample texts. However, in the case of Relevance, our method did not outperform the baselines in terms of Spearman correlation, suggesting that the dimension may need further refinement in checklist design.
5.3. Comparative analysis of five TA conditions (RQ3)
As this study differentiates the TA conditions, we observe how performance varies based on differences in attributes generated through the think-aloud method under five conditions: SL-TA, SH-TA, ML-TA, MH-TA, and Comb-TA. We then delve deeper into the effectiveness of combining the outcomes of humans’ and LLMs’ TA for generating checklists. The SL-TA indicates the correlations averaged across four individual LLM conditions, and the SH-TA represents the performance averaged over four different expert conditions.
5.3.1. Performance comparison across different TA conditions.
@@ 5.3. Comparative analysis of five TA conditions (RQ3)
Figure 4. Comparison of evaluation performance in Spearman correlations between single-TA and multiple-TA conditions across four dimensions.
@@
Figure 5. Comparison of two correlations measures (Spearman & Kendall-tau) over four dimensions-Coherence, Fluency, Consistency, and Relevance. The bars represent the correlations of different TA conditions: SL-TA (Avg), SH-TA (Avg), ML-TA, MH-TA, and Comb-TA. The performances of SL-TA (Avg) and SH-TA (Avg) represent the average values across the four different checklists created by each LLM and human expert.
As shown in Figure 4, using the attributes generated by multiple humans or LLMs results in better evaluation outcomes, including improved average performance, compared to using the ones thought by a single human or LLM, though the difference is not statistically significant (Table 4). Moreover, Figure 5 demonstrates that combining the TA outcomes from humans and LLMs for checklist construction (Comb-TA) outperforms all single-TA (SL-TA, SH-TA) and multiple-TA (ML-TA, MH-TA) based checklists, with the exception of the Kendall-tau score for Coherence. However, the gap is negligible and does not significantly affect the overall average performance. The improvement is significant in Fluency and Consistency. This suggests that certain aspects of text quality benefit more from this combination than others.
(a) Comparison of evaluation performance. As Rouge-L (X-axis) decreases, Spearman correlation (Y-axis) increases. The red line shows the upward trend across all dimensions. The decreasing Rouge-L signifies an increasing diversity of attributes within each condition.
(b) Topic distribution and standard deviation analysis using LDA.
Figure 6. Analysis of attributes and evaluation performance.
Figure 6(a) shows the correlation between Spearman values and self-ROUGE-L scores for attributes within the same TA condition. This suggests that checklists based on diverse attributes (i.e., with low similarity) enhance evaluation performance, supporting the idea that the multiple-TA condition helps create more fine-grained checklists. Furthermore, combining human and LLM outcomes leads to even more fine-grained checklists, resulting in the highest performance and diversity.
To examine potential biases within the attributes of each TA condition, we conducted latent dirichlet allocation (LDA) (Blei et al., 2003), as shown in Figure 6(b). This method generated five topics based on the attributes from the Comb-TA condition. We then analyzed these topics to determine whether the attributes from other TA condition were concentrated in specific topics, providing insights into potential biases. The results show that attributes from single-TA conditions, whether human or LLM, tend to cluster into a few topics, with some topics receiving no attributes at all, indicating bias. In contrast, attributes from the multiple-TA and Comb-TA conditions are more evenly distributed across topics. Additionally, the Comb-TA condition has the lowest standard deviation in topic distribution, suggesting that combining the TA outcomes of multiple humans and LLMs results in a less biased checklist.
@@
Table 4. Comparison of two correlations measures, Spearman ($\rho$) and Kendall-tau ($\tau$), over four dimensions-Coherence, Fluency, Consistency, and Relevance. The values are z-scores that indicate the statistical significance of the differences between the method pairs (L: LLM, H: Human, S: Single, M: Multiple, C: Combination), with asterisks (*) denoting the level of significance: * (p-vale <0.05), ** (p-value <0.01), and *** (p-value <0.001). The results are presented evaluator-wise.
Figure 7. Similarity of attributes in lexical and semantic perspectives. The bars represent each measurement of text similarity. The similarity of SL-TA (Avg) and SH-TA (Avg) represent the average values across the four different checklists created by each LLM and human expert. The lower the similarity, the more diverse the attributes are within a dimension.

### v2

@@ 4.4. Qualitative analysis
We conducted a focused group interview (FGI) (Morgan, 1996; Dilshad and Latif, 2013) with four human experts (E1, E2, E3, and E4) to qualitatively analyze the attributes produced by the humans (A) and LLMs (B) during the TA process. The experts compared both sets of attributes by responding to open-ended questions for each dimension (see Appendix F for the details of questions). These questions allowed us to gain deeper insights into the attributes and encouraged detailed expert feedback. Three of the experts had participated in the TA process, while one had not. Initially, we planned for all four TA participants to conduct the qualitative analysis, assuming their familiarity would lead to more refined insights. However, one expert with linguistics expertise withdrew for a personal reason. Hence, we invited another linguist, who had been also researching large language models, to participate. After collecting the experts’ responses through email, we identified keywords mentioned by the majority (at least three experts). We then filtered out answers containing the keywords. The results will be further discussed in section 5.3.2 and 5.3.3.
@@ 5. Results
5.1. Analysis of text evaluation performance (RQ1)
To examine the performance of InteractEval, we have conducted a series of experiments including a comparison against baseline text evaluation methods, examination of changes in model performance based on the number of components, and generalization assessment of model performance using a different dataset.
5.1.1. Performance comparison with baselines
We conducted a comparison against baseline evaluation methods to evaluate the performance of InteractEval. Building upon prior studies (Fu et al., 2023; Fabbri et al., 2021; Lee et al., 2024b; Liu et al., 2023), we analyzed the correlation between InteractEval and ground-truth scores using Spearman’s Rho, Kendall’s Tau coefficients, and the difference between predicted scores and human-labeled scores using mean absolute errors (MAE). For a fair comparison, we used the performance data of the seven non-LLM-based traditional methods (see Table 9), G-Eval (GPT-4), and CheckEval (GPT-3.5 Turbo and GPT-4), as reported in (Lee et al., 2024b), given that their experiments were conducted using the same experimental settings (sample-level correlation). Additionally, we reproduced the correlation results of G-Eval (GPT-3.5-Turbo) [^9] https://github.com/nlpyang/geval.git and the MAE results of the entire baselines at the sample level using the source code provided by the original authors (where available) and following the implementation details described in their papers.
Table 9. Sample-level performance of different dimensions on SummEval benchmark dataset measured by Spearman’s Rho ($\rho$), Kendall’s Tau ($\tau$), and Mean Absolute Error (m). The best results are highlighted in bold, and the second-best results are underlined.
@@ 5.1.1. Performance comparison with baselines
m ($\downarrow$) $\rho(\uparrow)$ $\tau(\uparrow)$ m ($\downarrow$) $\rho(\uparrow)$ $\tau(\uparrow)$ m ($\downarrow$) $\rho(\uparrow)$ $\tau(\uparrow)$ m ($\downarrow$) $\rho(\uparrow)$ $\tau(\uparrow)$ m ($\downarrow$) $\rho(\uparrow)$ $\tau(\uparrow)$
@@
ROUGE-L 2.008 0.172 0.122 3.305 0.240 0.167 3.372 0.317 0.235 2.356 0.420 0.301 2.760 0.287 0.206
BLEU 2.877 0.028 0.019 4.102 -0.079 -0.057 4.214 0.036 0.030 3.215 0.423 0.308 3.602 0.102 0.075
METEOR 1.632 0.009 0.008 2.679 0.011 0.008 2.813 0.101 0.074 1.847 0.275 0.182 2.243 0.099 0.068
BERTScore 1.095 -0.219 -0.151 0.715 -0.156 -0.113 0.642 -0.109 -0.076 1.601 0.462 0.476 1.013 -0.005 0.034
MOVERScore 0.978 0.340 0.252 1.866 0.345 0.247 1.961 0.384 0.277 1.050 -0.434 -0.315 1.464 0.159 0.115
BARTScore 1.150 0.475 0.343 1.617 0.395 0.285 1.838 0.457 0.341 1.148 0.602 0.455 1.438 0.482 0.356
UniEval 1.145 0.567 0.411 0.715 0.605 0.447 0.642 0.613 0.476 1.601 0.640 0.476 1.026 0.606 0.453
@@
G-Eval 1.351 0.412 0.393 2.750 0.354 0.333 0.871 0.338 0.322 0.844 0.361 0.321 1.454 0.366 0.342
CheckEval 1.135 0.440 0.347 1.078 0.393 0.327 1.694 0.600 0.515 0.851 0.305 0.252 1.190 0.435 0.360
InteractEval (1st) 1.185 0.583 0.536 1.860 0.630 0.604 1.558 0.734 0.702 0.998 0.614 0.576 1.400 0.640 0.605
InteractEval (2nd) 1.226 0.590 0.540 1.482 0.614 0.598 1.021 0.726 0.698 0.936 0.623 0.577 1.166 0.638 0.603
@@
G-Eval 1.084 0.619 0.470 2.195 0.629 0.482 1.216 0.664 0.517 1.399 0.617 0.464 1.474 0.632 0.483
CheckEval 0.944 0.573 0.428 0.934 0.632 0.493 1.429 0.706 0.611 0.684 0.570 0.438 0.998 0.620 0.493
InteractEval (1st) 0.847 0.649 0.603 0.356 0.799 0.769 1.042 0.783 0.761 0.919 0.626 0.587 0.791 0.714 0.680
InteractEval (2nd) 1.177 0.660 0.606 0.639 0.781 0.751 1.009 0.816 0.798 0.633 0.642 0.592 0.865 0.725 0.687
@@
Table 9 presents the main results, classifying the four dimensions into two key features by which generated texts are typically assessed: internal quality and external alignment (Zhu and Bhat, 2020; Wang et al., 2024c). Additionally, the table shows InteractEval’s performance in the first and second trials using both GPT-3.5-Turbo and GPT-4, demonstrating stable sample-level correlations. In the both trials, InteractEval with GPT-4 outperforms or is at least comparable to the traditional non-LLM-based metrics, such as ROUGE-L, BLEU, METEOR, BERTScore, MOVERScore, BARTScore, and UniEval, in terms of correlation coefficients. As noted in (Lee et al., 2024b), the results suggest that using a checklist enables an LLM to evaluate generated text in a manner more aligned with human judgment than traditional automatic metrics.
@@
When compared to G-Eval and CheckEval with the identical version of GPT, our model demonstrates superior performance. Moreover, even with GPT-3.5-Turbo, our framework outperforms other LLM-based baselines with GPT-4 or shows comparable performance, indicating its cost-saving potential while ensuring reasonable performance (Arefeen et al., 2024; Renze and Guven, 2024; Shekhar et al., 2024). These results highlight the broad capability of our proposed framework, showing that using TA and the checklist constructed through the combined TA results of humans and LLMs lead to accurate and cost-effective evaluations across the four text quality dimensions of text summarization.
In terms of MAE performance, there is no clear consistent trend, although InteractEval with GPT-4 achieves the best performance on average. Specifically, the results show that models with high correlations to human raters do not necessarily produce lower MAE errors. This discrepancy arises from the theoretical differences between MAE, which measures the average absolute difference between predicted and actual scores (Willmott and Matsuura, 2005), and correlation, which evaluates the strength and direction of a linear relationship between the scores (Lee Rodgers and Nicewander, 1988). Unlike correlation, MAE does not account for the direction of errors and cannot distinguish between overestimations and underestimations. Taken together, the sample-level analysis conducted in relation to human assessment scores as ground-truth indicate that correlation analysis on the basis of Spearman’s Rho or Kendall’s Tau are more sensitive than the analysis conducted with Mean Absolute Error.
Figure 6. Performance of InteractEval across Different Component Numbers: Two correlation measures (y-axis) are compared across the four evaluation dimensions—Coherence, Fluency, Consistency, and Relevance—by varying the number of components (x-axis) extracted from Comb-TA attributes. The reported correlation values represent the average results from the first and second trials.
5.1.2. Performance comparison with component numbers.
As previously mentioned, we prompted an LLM to extract five components from TA attributes before generating questions and constructing checklists. To evaluate the impact of the number of components on InteractEval’s performance, we conducted an ablation study, varying the number of components from three to nine, as shown in Figure 6. The results indicate that our model achieves optimal performance when using checklists constructed with five components derived from TA attributes. We believe these performance differences are reflective of the difficulty of capturing sufficient key sub-concepts related to each evaluation dimension when extremely few components are extracted, and the redundancy that arises when too many components are extracted. Therefore, we empirically confirm through experiments that extracting five components strikes an optimal balance, enabling the identification of the most relevant key sub-concepts for all dimensions and facilitating the creation of the most effective checklists.
Figure 7. Performance Comparison on ELLIPSE dataset: Sample-level correlations of different dimensions on ELLIPSE benchmark dataset are measured by Spearman’s Rho ($\rho$) and Kendall’s Tau ($\tau$). The reported results represent the average performance across the first and second trials (see Appendix E for exact values).
Table 10. Sample-level correlations of different dimensions on SummEval benchmark dataset are measured using Spearman’s Rho ($\rho$) and Kendall’s Tau ($\tau$). Entire models utilized GPT-4 as the evaluator. The model names in parentheses in InteractEval refer to the models used for the TA process. The best results are highlighted in bold, and the second-best results are underlined.
@@
5.1.3. Evaluation performance on ELLIPSE dataset
To provide supplementary evaluation beyond the text summary benchmark dataset and examine the generalizability of InteractEval across different tasks, we employed our method using the ELLIPSE dataset. We followed the same procedures as the SummEval dataset including attribute collection through think aloud, checklist construction, and checklist-based evaluation. Additionally, we conducted the human TA procedure with four linguistic experts, considering that the dataset is related to English language education (see appendix C for detailed information of the participants). None of the baselines report their performance on the dataset; accordingly, the baseline results are based on our implementation following the experimental settings outlined in their original papers or source codes (if available). Note that, in CheckEval (Lee et al., 2024b), authors filtered LLM-generated checklists in the final stage; therefore, we followed the same process after obtaining the checklists. As with the SummEval dataset, we sampled 10% of the original dataset twice and ran InteractEval and the baselines on the two samples (referred to as the first trial and second trial below).
Figure 7 presents the average performance of the top four models across the first and second trials. The results show that InteractEval surpasses all baselines across entire dimensions, demonstrating the broader applicability of InteractEval on text evaluation. However, unlike with the SummEval dataset, our method using GPT-3.5-Turbo does not consistently outperform other baselines, suggesting that the cost-effectiveness of InteractEval depends on the specific dataset used. Because InteractEval produces the best performance outcomes when using GPT-4 as the evaluator, the remaining experiments were all conducted with GPT-4.
5.2. Analysis of TA effectiveness (RQ2)
@@
The proposed framework initially collects various attributes from humans and LLMs that should be considered while designing checklists per dimension. Unlike prior research, we utilize TA to encourage LLMs to generate a wider range of text attributes, which form the basis for checklists across different evaluation dimensions. To examine how the TA process affects the evaluation performance of our framework, we compare its performance when using checklists constructed from a single LLM’s attributes under the TA process against other LLM-based baselines on the SummEval dataset. As shown in Table 10, the performance of InteractEval across the four evaluation dimensions varies depending on the LLM used for the TA process. For each dimension, at least one of the four InteractEval conditions with a single-LLM TA surpasses the baselines, G-Eval (Liu et al., 2023) and CheckEval (Lee et al., 2024b), except for the Spearman’s Rho correlation of Relevance. The results show that incorporating TA improves the generation of text attributes by LLMs to some extent. In particular, our method, using single-LLM-TA outcomes with GPT-4, outperforms other LLM-based baselines, except in Spearman’s Rho correlation for Relevance. For overall average, irrespective of whether the correlation is assessed via Spearman’s Rho or Kendall’s Tau, InteractEval with single-LLM TA (GPT-4) surpasses all of the baselines, without exception. These results suggest that even with just one LLM, particulary with GPT-4, the TA process leads to more refined checklists, improving evaluation outcomes. By encouraging divergent analysis (Ericsson and Simon, 1980; Ericsson, 6; Lewis, 1982), the TA process enables LLMs to capture more detailed aspects of the dimension rubrics based on sample texts. In the end, the study results show that combining the TA outcomes from the four LLMs complements each other, contributing to removing idiosyncratic differences and overcoming each one’s weaknesses, and consequently results in checklists of higher quality.
@@
Figure 8. Performance Comparison Between Single-TA and Multiple-TA Conditions: Evaluation performance in Spearman’s Rho ($\rho$) correlations is compared between single-TA and multiple-TA conditions across the four dimensions. Single-Human and Single-LLM represent the average performance results from four human participants and four LLMs, respectively.
@@
Figure 9. Comparison of two correlation measures (Spearman’s Rho ($\rho$) & Kendall’s Tau ($\tau$)) across the four evaluation dimensions - Coherence, Fluency, Consistency, and Relevance. The bars represent the correlations of different TA conditions: SL-TA (Avg), SH-TA (Avg), ML-TA, MH-TA, and Comb-TA. The performance of SL-TA (Avg) and SH-TA (Avg) is based on the average values across the four different checklists created by each LLM and human expert, respectively.
Table 11. Comparison of two correlation measures, Spearman’s Rho ($\rho$) and Kendall’s Tau ($\tau$), over four dimensions-Coherence, Fluency, Consistency, and Relevance. The values are z-scores that indicate the statistical significance of the differences between the method pairs (L: LLM, H: Human, S: Single, M: Multiple, C: Combination), with asterisks (*) denoting the level of significance: * (p-vale <0.05), ** (p-value <0.01), and *** (p-value <0.001).
@@
5.3. Comparative analysis of five TA conditions (RQ3)
As this study differentiates the TA conditions, we observe how performance varies based on differences in attributes generated through the TA method under the five conditions: SL-TA, SH-TA, ML-TA, MH-TA, and Comb-TA. We then delve deeper into the effectiveness of combining the outcomes of humans’ and LLMs’ TA outcomes for generating checklists. The SL-TA score represents the correlation coefficients averaged across the four individual LLM conditions, and the SH-TA score indicates the coefficients averaged across the four different human expert conditions on the SummEval dataset.
5.3.1. Performance comparison across different TA conditions.
As shown in Figure 8, using the attributes generated by multiple humans or multiple LLMs results in better evaluation outcomes, including improved average performance, compared to using the ones produced by a single human or a single LLM, though the difference is not statistically significant (Table 11). Moreover, Figure 9 demonstrates that combining the TA outcomes from humans and LLMs for checklist construction (Comb-TA) outperforms all single-TA (SL-TA, SH-TA) and all multiple-TA (ML-TA, MH-TA) based checklists. While there is a minor exception in the Kendall’s Tau score for Coherence, the gap is negligible and does not significantly affect the overall average performance. The improvement is significant in Fluency and Consistency (Table 11), indicating that certain aspects of text quality benefit more from this combination than others.
(a) Comparison of evaluation performance. As Rouge-L (X-axis) decreases, Spearman’s Rho ($\rho$) correlation (Y-axis) increases. The red line shows the upward trend across all dimensions. The decreasing Rouge-L signifies an increasing diversity of attributes within each condition.
(b) Topic Distribution and Standard Deviation Analysis using LDA: The upper figure shows that attributes from single-TA conditions tend to cluster around a few topics, whereas attributes from multiple-TA conditions are more evenly distributed across various topics. The lower figure indicates that the standard deviation of Comb-TA is the lowest, suggesting reduced bias.
Figure 10. Analysis of Attributes and Evaluation Performance.
Figure 10(a) shows the correlation coefficients between Spearman’s Rho ($\rho$) values and self-ROUGE-L scores for attributes within the same TA condition. The results show that those checklists based on diverse attributes (i.e., with low similarity) enhance evaluation performance, supporting the idea that the multiple-TA condition helps create more fine-grained checklists and enhance the evaluation performance. Furthermore, combining human and LLM outcomes leads to even more fine-grained checklists, resulting in the highest performance and diversity.
To examine potential biases, particularly short-sighted or myopic cognitive biases, within the attributes of each TA condition, we conducted latent dirichlet allocation (LDA) (Blei et al., 2003), as shown in Figure 10(b). This analysis generated five topics based on the attributes from the Comb-TA condition. The threshold of five is consistent with the maximum number of components extracted from each of the TA conditions. We then analyzed how the attributes from the TA conditions were concentrated in specific topics. As shown in Figure 10(b), the results show that attributes from the single-TA conditions, whether human or LLM, tend to cluster into a fewer topics, with some topics receiving no attributes at all, indicating the existence of myopic bias. In contrast, attributes from the multiple-TA and Comb-TA conditions are more evenly distributed across these five topics. Additionally, the Comb-TA condition has the lowest standard deviation in topic distribution, suggesting that combining the TA outcomes of multiple humans and LLMs results in a less biased checklist.
Figure 11. Similarity of Attributes in Lexical and Semantic Perspectives: The bars represent the measurement of text similarity. The similarity of SL-TA (Avg) and SH-TA (Avg) represents the average values across the four different checklists created by each LLM and each human expert, respectively. The lower the similarity, the more diverse the attributes are within a dimension.